(1). Let f: Z → Z be given by f(x) = x2. Find F-1(D) where (a) D = {2,4,6,8, 10, 12, 14, 16}. (b) D={-9, -4,0, 16, 25}. (c) DQuestion4 please
Answer #1

(4)

Bijection : A function is bijection if it is both one - one and onto.

One -one : Every element in the range set of function has a unique pre image in domain set.

Onto: Every element in the co-domain set has atleast one pre image in the domain set, i.e. range set = co-domain set.

We prove these properties for the chosen function f(x).

And the find its inverse.

I (4.) So find: Bijection from W-N w = set of whole numbers Solutions - fake ((x) = x+ 1 N = set of natural = numbers + XEW= y = f(x) =) y = S(+1 = x=yt i f (y) = x=yt # yein Let ftly)= gly) then inverse of f =) : INJW Scanned with by Igly) = y1)